OpenStudy (anonymous):

Which polynomial is prime? x3 + 49 x3 – 1000 x3 + 1 x3 – 64

OpenStudy (anonymous):

I think its C

jimthompson5910 (jim_thompson5910):

actually you can use the sum of cubes factoring rule to factor x^3 + 1 to get (x+1)(x^2 - x + 1)

jimthompson5910 (jim_thompson5910):

sum of cubes factoring rule a^3 + b^3 = (a+b)(a^2 - ab + b^2)

jimthompson5910 (jim_thompson5910):

in the case of x^3 + 1= (x+1)(x^2 - x + 1), we have a = x and b = 1

OpenStudy (anonymous):

So am I correct? If not dont tell me ill figure it out

OpenStudy (mayankdevnani):

nope, you are wrong !

OpenStudy (mayankdevnani):

the polynomial is prime which can't be factored out

jimthompson5910 (jim_thompson5910):

@hippo211 since i was able to factor x^3 + 1 to get (x+1)(x^2 - x + 1), that means x^3+1 is NOT prime you only say something is prime if you cannot factor

OpenStudy (mayankdevnani):

correct ^^^

OpenStudy (mayankdevnani):

@hippo211 any problem ?

OpenStudy (anonymous):

thank you